OpenStudy (anonymous):

In horse racing, a trifecta is a bet for which you must pick the first, second, and third place winners in their proper order in order to win a payoff. If there are fourteen horses racing, and you randomly select three as your bet in the trifecta, what is the probability that you will win?

OpenStudy (anonymous):

you only pick one right, so it is one out of the number of ways the the first 3 horses can finish, which is \(14\times 13\times 12\)

OpenStudy (anonymous):

\[\frac{1}{14\times 13\times 12}\] in other words

OpenStudy (anonymous):

thank you!!! ive called like 6 people and no one could figure it out!